Starting with a single bacterium you would expect to have bacteria after 4 generations.
Answer #1

Starting with a single bacterium you would expect to have 16 bacteria after 4 generations

Explanation: Bacteria reproduce via binary fission. This process produces two bacteria from one parent cell. A single bacterium (say zeroth generation) produces two bacteria after one generation, which can then produce four bacteria in the second generation, which, in turn, can produce eight bacteria in the third generation and in the fourth generation this 8 bacterium would divide to produce 16 bacterium. Ideally given by 2^n = 2^4 = 16.